Grade 11Algebra

For what values of a belongs to R, do the roots of x2-2x-(a2-1)=0, lie between the roots of x2-2(a+1)x+a(a-1)=0,
Could you pls assist me with detailed step by step solution.

Surya
8 Years agoGrade 11

1 Answer

Surya
8 Years ago
f(x)=0
=x2-2x-(a2-1)=0
=(x-1)2-a2=0
So, x=1\pma
Let \alpha=1-a and Let \beta=1+a
Now lets Consider 3 cases:
1)\Delta>0= a\in(-1/3,Infinity)----(i)
2)g(\alpha)<0=a\in(-1/4,1)------(ii)
3)g(\beta)<0=a\in(-1/3, Infinity)---------(iii)
Now taking intersection of all three cases, we get,
a\in(-1/4,1)